OpenStudy (anonymous):

A truck covers 40.0 m in 8.50 s while smoothly slowing down to a final velocity of 2.80 m/s. (a) Find the truck's original speed. (b) Find acceleration

OpenStudy (anonymous):

Here s =40.0 m t = 8.50 s v = 2.80 m/s (a) original speed = initial speed = u \[s = \frac{ u+v }{ 2 } \times t\] (b) \[a = \frac{ v-u }{ t }\]